OpenStudy (anonymous):

please help Rewrite 2cos 75°sin 75° using a double-angle identity

OpenStudy (jdoe0001):

http://www.efunda.com/math/trig_functions/images/double.gif <--- try using the sine identity

OpenStudy (anonymous):

sin 2x=2sin x cos x

OpenStudy (anonymous):

so sin 2(75)= 2sin (75) cos (75)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

sin150=sin(180-30)=?

OpenStudy (anonymous):

\[\sin \left( 180-\theta \right)=\sin \theta\]

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

yw